# JEL Classifications: C61; D15; D52; E21; G51

### v2

@@ 2.2. Optimality: Definitions and Fundamental Properties
For all $z\in\mathsf{Z}$, $u(\cdot,z)$ is twice differentiable on $(0,\infty)$ with $u^{\prime}(\cdot,z)>0$, $u^{\prime\prime}(\cdot,z)<0$, $u^{\prime}(c,z)\to\infty$ as $c\to 0$, and $u^{\prime}(c,z)<1$ as $c\to\infty$.
Since we did not assume that the utility function is bounded or the discount factor $\beta_{t}$ is less than one, the lifetime utility (1) may not be well defined. To overcome this issue, we define optimality by the overtaking criterion of Brock (1970). Given a feasible policy $c$, we consider the expected sum of utilities over a finite horizon up to time $T$, denoted by
@@
$u^{\prime}\left(c(w,z),z\right)=\max\left\{\mathbbm{E}_{z}\hat{\beta}\hat{R}u^{\prime}\big(c(\hat{w},\hat{Z}),\hat{Z}\big),u^{\prime}(w,z)\right\}$ (5)
@@
$\lim_{t\to\infty}\mathbbm{E}_{w,z}\left(\prod_{i=0}^{t}\beta_{i}\right)u^{\prime}\left(c(w_{t},Z_{t}),Z_{t}\right)w_{t}=0.$ (6)
@@ 2.3. Existence and Computability of Optimal Consumption
$u^{\prime}\left(\xi,z\right)=\max\left\{\mathbbm{E}_{z}\hat{\beta}\hat{R}u^{\prime}\left(c\big(\hat{R}(w-\xi)+\hat{Y},\hat{Z}\big),\hat{Z}\right),u^{\prime}\big(w,z\big)\right\}.$ (8)
@@
$r(A):=\max\{|\alpha|:\alpha\text{ is an eigenvalue of }A\}.$
@@ 2.4. Fundamental Properties of Consumption and Saving
[…]
@@
3. Asymptotic Properties of Consumption
In this section, we study the asymptotic properties of the consumption function when there are risk aversion shocks. Throughout we assume that Assumptions 2.12.2 hold.
@@ 3.1. Setups
$u(c,z)=\begin{cases}\frac{c^{1-\gamma(z)}}{1-\gamma(z)},&\text{if }\gamma(z)>0\text{ and }\gamma(z)\neq 1,\\ \log c,&\text{if }\gamma(z)=1,\end{cases}$ (10)
@@
Assume that the Markov process $\{Z_{t}\}$ can be decomposed as $Z_{t}=(\bar{Z}_{t},\tilde{Z}_{t})$, where $\{\bar{Z}_{t}\}$ and $\{\tilde{Z}_{t}\}$ are independent Markov processes taking values in $\bar{\mathsf{Z}}:=\{\bar{z}_{1},\dots,\bar{z}_{N}\}$ and $\tilde{\mathsf{Z}}:=\{\tilde{z}_{1},\dots,\tilde{z}_{M}\}$, with transition matrices $\bar{P}=(\bar{p}_{ij})_{1\leq i,j\leq N}$ and $\tilde{P}=(\tilde{p}_{ij})_{1\leq i,j\leq M}$, respectively. In this setting, the transition probability of $\{Z_{t}\}$ satisfies $P=\bar{P}\otimes\tilde{P}$, where $\otimes$ denotes the Kronecker product. In addition, assume that the level of relative risk aversion is driven solely by $\{\bar{Z}_{t}\}$ and satisfies
@@ 3.2. The Asymptotic MPCs
$\lim_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}=0\quad\text{for all }\,j=1,\dots,M.$ (12)
@@
$\lim_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}=0\quad\text{for all }z_{ij}\in\mathsf{Z}\text{ with }i\neq 1.$ (13)
@@
$(F_{i}x)(\tilde{z}_{j}):=\left(1+(G_{i}x)(\tilde{z}_{j})^{1/\gamma_{i}}\right)^{\gamma_{i}},\qquad j\in\{1,\dots,M\}.$ (15)
@@
$\lim_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}=x_{i}^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}>0\quad\text{for all }\,j=1,\dots,M.$ (16)
@@
Theorems 3.33.4 address the complementary cases of Theorem 3.1, corresponding to $\sum_{\ell=1}^{i-1}\bar{p}_{\ell i}=0$ with $\bar{p}_{ii}=1$ and $\bar{p}_{ii}=0$, respectively. Our final result covers the remaining case with $\sum_{\ell=1}^{i-1}\bar{p}_{\ell i}=0$ and $\bar{p}_{ii}\in(0,1)$. Together, these results exhaust all possible configurations of the transition probability $\bar{P}$.
@@
$\bar{c}(z):=\lim_{w\to\infty}\frac{c^{*}(w,z)}{w}.$
Example 3.1 (Lower triangular transition matrix).
@@ 3.3. The Asymptotic Saving Rate
Based on (2) and some simple algebra, we can show that [^10] See Section 2.5 of Ma and Toda (2021). Here, we define $x^{+}=\max\{x,0\}$ and $x^{-}=-\min\{x,0\}$ for each $x\in\mathbbm{R}$.
@@
When the rate of return on wealth is positive (i.e., $R_{t}>1$) and the asymptotic MPC is zero, the asymptotic saving rate converges to 100%. Conversely, if the asymptotic MPC is positive and sufficiently large, the asymptotic saving rate becomes negative, indicating that the precautionary saving motive vanishes as wealth grows without bound.
@@ 4. Quantitative Analysis
The preceding theoretical analysis shows that allowing for dynamically evolving risk aversion fundamentally alters the consumption–saving dynamics of the model. In this section, we complement the theory with a quantitative analysis that serves two purposes. First, using a parsimonious quantitative specification, we provide numerical illustrations that render the main theoretical results transparent and economically intuitive. Second, within a more general quantitative environment, we compute generalized impulse response functions to elucidate the economic mechanisms through which preference shocks generate qualitative changes in model dynamics.
@@ 4.1. Model Specification
For our quantitative analysis, we consider an optimal savings problem with shocks to risk aversion. The discount factor and the gross rate of return on wealth are held constant over time, with $\beta_{t}\equiv\beta$ and $R_{t}\equiv R$. Risk aversion follows a discretized autoregressive process of order one:
@@
This specification is a special case of our theoretical framework, with the preference state given by $\bar{Z}_{t}=\gamma_{t}$. For the nonfinancial income process, we allow for alternative specifications depending on the quantitative experiment under consideration. The model is solved using the endogenous grid method of Carroll (2006). Below we describe the key elements of the quantitative setup. Detailed implementation and algorithmic steps are provided in Appendix E.
In our quantitative experiments, we work at an annual frequency and set $\beta=0.95$ and $R=1.02$. In our baseline setting, the risk aversion process is parameterized by $\rho_{\gamma}=0.6$, $\mu_{\gamma}=4$, and $\sigma_{\gamma}=1$, and discretized into a finite-state Markov chain using the method of Tauchen (1986). The number of discrete states is allowed to vary across experiments. In particular, for the impulse response analysis in Section 4.3, we employ a relatively fine discretization to improve approximation accuracy. Across all specifications, the minimum, median (and mean), and maximum values of $\gamma_{t}$ are $1$, 44, and $7$, respectively, which lie well within the range commonly used in empirical research.
@@ 4.2. The Asymptotic Saving Rates
To ensure a fair comparison, we contrast the optimal consumption function when $\gamma_{t}$ is fixed at its stationary mean $\mathbbm{E}\gamma_{t}=4$ in our stochastic risk aversion model with that of the standard optimal savings problem featuring CRRA utility with a constant risk aversion coefficient $\gamma=4$. Figure 1 plots the results. Despite the fact that current risk aversion is identical across the two models and that there is a high probability of remaining at the same level of risk aversion next period, consumption is substantially lower under stochastic risk aversion (red dotted line) than under constant risk aversion (blue solid line), reflecting the precautionary behavior induced by the possibility of future changes in risk aversion.
@@
The right panel plots the corresponding saving rates. In the standard optimal savings model, saving rates are negative and converge to their theoretical asymptotic value of $40.31\%$, as given by (17), in stark contradiction with the empirical evidence on wealthy households. In contrast, under stochastic risk aversion, the saving rate converges to its theoretical limit of $100\%$. Taken together, these results illustrate that even small probabilities of shifts in risk aversion can have sizable effects on the saving behavior, suggesting that stochastic risk aversion offers a promising mechanism for generating persistently high saving rates among wealthy households, in line with the patterns observed in the data.
@@ 4.3. Impulse Response Analysis
To ensure sufficient numerical accuracy, we discretize $\{\gamma_{t}\}$ into a 121-state Markov chain following Tauchen (1986). The persistent income component $\{\eta_{t}\}$ is discretized into 5 states, while expectations over the transitory shock $\{\nu_{t}\}$ are computed using a 7-point Gauss–Hermite quadrature. Sensitivity analysis confirms that the results are robust to these parametric choices.
@@
The top row of Figure 3 illustrates heterogeneity with respect to initial wealth, holding risk aversion fixed at its stationary mean $\gamma_{t-1}=4$. The left panel shows that an increase in risk aversion induces an immediate and persistent decline in consumption, with the magnitude of the response rising sharply in wealth. When wealth rises from $w_{t-1}=5$ to $10$ and $50$, contemporaneous consumption falls by 4.67%, 7.89%, and 20.23%, respectively, with effects lasting roughly 2–6 years. This nonlinearity reflects the fact that low-wealth households are constrained in their ability to adjust savings, so consumption responds weakly to preference shocks, whereas wealthy households can reallocate intertemporal resources more flexibly.
@@
The middle row of Figure 3 illustrates heterogeneity with respect to initial risk aversion, holding initial wealth fixed at $w_{t-1}=10$. When initial risk aversion is high ($\gamma_{t-1}=6$), the same shock generates relatively modest responses: consumption falls by 4.84% and the saving rate rises by 5.53%. By contrast, when agents start from a low risk aversion state ($\gamma_{t-1}=1$), consumption drops sharply by 19.26% and saving increases by 32.21%. This asymmetry reflects state-dependent amplification: a positive shock represents a larger proportional change in effective prudence when baseline risk aversion is lower, triggering a stronger re-optimization of intertemporal choices. Consistent with this mechanism, the decline in consumption volatility is also largest at low initial risk aversion.
@@
5. Conclusion
We developed a generalized optimal savings model with preference shocks, allowing for time-varying and state-dependent risk aversion. This extension fundamentally reshapes the asymptotic behavior of consumption and saving. Our analysis shows that zero asymptotic MPCs arise naturally whenever agents face a positive probability of becoming less risk averse, even in the absence of return risk or stochastic discounting.
Quantitative illustrations further confirm that this mechanism can generate large and persistent variations in consumption and saving across states, consistent with observed wealth accumulation patterns. Taken together, our framework provides a theoretically robust and empirically consistent explanation for the persistent accumulation of wealth and the high saving rates observed among the rich.
By linking saving behavior to the dynamics of risk preferences, our model highlights a potential channel through which heterogeneous risk attitudes may influence aggregate saving and offers insights relevant for understanding wealth inequality and asset market outcomes. A detailed exploration of these implications is left for future research.
@@ Appendix A Preliminaries
$(L_{\varphi}^{n}\,h)(z)=\mathbbm{E}_{z}\prod_{t=1}^{n}\varphi_{t}h(Z_{t}),$ (19)
where $L_{\varphi}^{n}$ is the $n$-th composition of the operator $L_{\varphi}$ with itself, or equivalently, the $n$-th power of the matrix $L_{\varphi}$.
@@
Proof.
The nonnegativity of $L_{\varphi}$ and Theorem 9.1 of Krasnosel’skii et al. (2012) imply that $r(L_{\varphi})=\lim_{n\to\infty}\|L_{\varphi}^{n}\,h\|^{1/n}$, where $\|\cdot\|$ is any norm on $\mathbbm{R}^{\mathsf{Z}}$ and $h$ is a strictly positive function on $\mathsf{Z}$. For $h\equiv 1$ and $\|f\|=\max_{z\in\mathsf{Z}}|f(z)|$, we have
$r(L_{\varphi})=\lim_{n\to\infty}\left(\max_{z\in\mathsf{Z}}\,L_{\varphi}^{n}\,\mathbbm{1}(z)\right)^{1/n}=\lim_{n\to\infty}\left(\max_{z\in\mathsf{Z}}\mathbbm{E}_{z}\prod_{t=1}^{n}\varphi_{t}\right)^{1/n}.$ (20)
Since $r(L_{\varphi})<1$, it follows by the definition of the limit that the stated claim holds. This completes the proof. ∎

### v3

@@ 2.2. Optimality: Definitions and Fundamental Properties
For all $z\in\mathsf{Z}$, $u(\cdot,z)$ is twice differentiable on $(0,\infty)$ with $u^{\prime}(\cdot,z)>0$, $u^{\prime\prime}(\cdot,z)<0$, and $u^{\prime}(c,z)\to\infty$ as $c\to 0$.
Since we did not assume that the utility function is bounded or the discount factor $\beta_{t}$ is less than one, the lifetime utility (1) may not be well defined in $\mathbbm{R}$. To overcome this issue, we define optimality by the overtaking criterion of Brock (1970). Given a feasible policy $c$, we consider the expected sum of utilities over a finite horizon up to time $T$, denoted by
@@
(5) $u^{\prime}\left(c(w,z),z\right)=\max\left\{\mathbbm{E}_{z}\hat{\beta}\hat{R}u^{\prime}\big(c(\hat{w},\hat{Z}),\hat{Z}\big),u^{\prime}(w,z)\right\}$
@@
(6) $\lim_{t\to\infty}\mathbbm{E}_{w,z}\left(\prod_{i=0}^{t}\beta_{i}\right)u^{\prime}\left(c(w_{t},Z_{t}),Z_{t}\right)w_{t}=0.$
@@ 2.3. Existence and Computability of Optimal Consumption
(8) $u^{\prime}\left(\xi,z\right)=\max\left\{\mathbbm{E}_{z}\hat{\beta}\hat{R}u^{\prime}\left(c\big(\hat{R}(w-\xi)+\hat{Y},\hat{Z}\big),\hat{Z}\right),u^{\prime}\big(w,z\big)\right\}.$
@@
$r(A):=\max\{|\lambda|:\lambda\text{ is an eigenvalue of }A\}.$
@@ 2.4. Fundamental Properties of Consumption and Saving
If $\{Y_{1t}\}$ and $\{Y_{2t}\}$ are two income processes satisfying $Y_{1t}\leq Y_{2t}$ for all $t$ and $c_{1}^{*}$ and $c_{2}^{*}$ are the corresponding optimal consumption functions, then $c_{1}^{*}\leq c_{2}^{*}$ pointwise on $S\SS$.
@@
3. Asymptotic Properties of Consumption and Saving
In this section, we study the asymptotic properties of consumption and saving when there are risk aversion shocks. Throughout we assume that Assumptions 2.12.2 hold.
@@ 3.1. Setups
(10) $u(c,z)=\begin{cases}\frac{c^{1-\gamma(z)}}{1-\gamma(z)},&\text{if }\gamma(z)>0\text{ and }\gamma(z)\neq 1,\\ \log c,&\text{if }\gamma(z)=1,\end{cases}$
@@
Assume that the Markov process $\{Z_{t}\}$ can be decomposed as $Z_{t}=(\bar{Z}_{t},\tilde{Z}_{t})$, where $\{\bar{Z}_{t}\}$ and $\{\tilde{Z}_{t}\}$ are independent Markov processes taking values in $\bar{\mathsf{Z}}:=\{\bar{z}_{1},\dots,\bar{z}_{N}\}$ and $\tilde{\mathsf{Z}}:=\{\tilde{z}_{1},\dots,\tilde{z}_{M}\}$, with transition matrices $\bar{P}=(\bar{p}_{ii^{\prime}})_{1\leq i,i^{\prime}\leq N}$ and $\tilde{P}=(\tilde{p}_{jj^{\prime}})_{1\leq j,j^{\prime}\leq M}$, respectively. In this setting, the transition probability of $\{Z_{t}\}$ satisfies $P=\bar{P}\otimes\tilde{P}$, where $\otimes$ denotes the Kronecker product. In addition, assume that the level of relative risk aversion is driven solely by $\{\bar{Z}_{t}\}$ and satisfies
@@ 3.2. The Asymptotic MPCs
(12) $\lim_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}=0\quad\text{for all }\,j=1,\dots,M.$
@@
(13) $\lim_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}=0\quad\text{for all }z_{ij}\in\mathsf{Z}\text{ with }i\neq 1.$
@@
(15) $(F_{i}x)(\tilde{z}_{j}):=\left(1+(G_{i}x)(\tilde{z}_{j})^{1/\gamma_{i}}\right)^{\gamma_{i}},\qquad j\in\{1,\dots,M\}.$
@@
(16) $\lim_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}=x_{i}^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}>0\quad\text{for all }\,j=1,\dots,M.$
@@
Theorems 3.33.4 address the complementary cases of Theorem 3.1, corresponding to $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}=0$ with $\bar{p}_{ii}=1$ and $\bar{p}_{ii}=0$, respectively. Our final result covers the remaining case with $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}=0$ and $\bar{p}_{ii}\in(0,1)$. Together, these results exhaust all possible configurations of the transition probability $\bar{P}$.
@@
$\bar{c}(z):=\lim_{w\to\infty}\frac{c^{*}(w,z)}{w}.$
Example 3.1 (Lower triangular transition matrix).
@@ 3.3. The Asymptotic Saving Rate
Based on (2) and some simple algebra, we can show that [^9] Here, we define $x^{+}=\max\{x,0\}$ and $x^{-}=-\min\{x,0\}$ for each $x\in\mathbbm{R}$.
@@
When the rate of return on wealth is positive (i.e., $R_{t}>1$) and the asymptotic MPC is zero, the asymptotic saving rate converges to 100%. Conversely, if the asymptotic MPC is positive and sufficiently large, the asymptotic saving rate becomes negative, implying that precautionary saving motive vanishes as wealth grows without bound.
@@ 4. Quantitative Analysis
In this section, we complement the theory with a quantitative analysis that serves two purposes. First, using a parsimonious quantitative specification, we provide numerical illustrations that render the main theoretical results transparent and economically intuitive. Second, within a more general quantitative environment, we compute generalized impulse response functions to elucidate the economic mechanisms through which preference shocks generate qualitative changes in model dynamics.
@@ 4.1. Model Specification
Consider an optimal savings problem with risk aversion shocks. The discount factor and the gross rate of return on wealth are held constant over time with $\beta_{t}\equiv\beta$ and $R_{t}\equiv R$. Risk aversion follows a discretized AR(1) process:
@@
This specification is a special case of our theory, with the preference state given by $\bar{Z}_{t}=\gamma_{t}$. For nonfinancial income, we allow for alternative specifications depending on the quantitative experiment under consideration. The model is solved using the endogenous grid method of Carroll (2006). Detailed implementation and algorithmic steps are provided in Appendix E.
In our quantitative experiments, we work at an annual frequency and set $\beta=0.95$ and $R=1.02$. In our baseline setting, risk aversion is parameterized by $\rho_{\gamma}=0.6$, $\mu_{\gamma}=4$, and $\sigma_{\gamma}=1$, and discretized into a finite-state Markov chain using the method of Tauchen (1986). The number of discrete states is allowed to vary across experiments. In particular, for the impulse response analysis in Section 4.3, we employ a fine discretization to improve approximation accuracy. In all specifications, the minimum, median (and mean), and maximum values of $\gamma_{t}$ are $1$, 44, and $7$, respectively, which lie well within the range commonly used in empirical research.
@@ 4.2. The Asymptotic Saving Rates
To ensure a fair comparison, we contrast the optimal consumption function when $\gamma_{t}$ is fixed at its stationary mean $\mathbbm{E}\gamma_{t}=4$ in our stochastic risk aversion model with that of the standard optimal savings problem featuring CRRA utility with a constant risk aversion coefficient $\gamma=4$. Figure 1 plots the results. [^10] The main difference between the left and right panels lies in the range of wealth displayed. Moreover, because the right panel is set on a log scale, the consumption function appears nonconcave, though it is in fact concave. A log scale is similarly used in Figure 2 when required. Despite the fact that current risk aversion is identical across the two models and that there is a high probability of remaining at the same level of risk aversion next period, consumption is substantially lower under stochastic risk aversion (red dotted line) than under constant risk aversion (blue solid line), reflecting the precautionary behavior induced by the possibility of future changes in risk aversion.
@@
The right panel plots the corresponding saving rates. In the standard optimal savings model, saving rates are negative and converge to their theoretical asymptotic value of $40.31\%$, as given by (17), in stark contradiction with the empirical evidence on wealthy households. In contrast, under stochastic risk aversion, the saving rate converges to its theoretical limit of $100\%$. Taken together, these results illustrate that even minor probabilities of shifts in risk aversion can have sizable effects on the saving behavior, suggesting that stochastic risk aversion offers a promising mechanism for generating persistently high saving rates among wealthy households, in line with the patterns observed in the data.
@@ 4.3. Impulse Response Analysis
To ensure sufficient numerical accuracy, we discretize $\{\gamma_{t}\}$ into a 121-state Markov chain following Tauchen (1986). In addition, the persistent income component $\{\eta_{t}\}$ is discretized into 5 states, while expectations over the transitory shock $\{\nu_{t}\}$ are computed using a 7-point Gauss–Hermite quadrature. We also conducted experiments using the discretization method of Rouwenhorst (1995). Sensitivity analysis confirms that the results are robust to both parametric choices and discretization methods.
@@
The top row of Figure 3 illustrates heterogeneity with respect to initial wealth, holding risk aversion fixed at its stationary mean $\gamma_{t-1}=4$. The top-left panel shows that an increase in risk aversion induces an immediate and persistent decline in consumption, with the magnitude of the response rising sharply in wealth. When wealth rises from $w_{t-1}=5$ to $10$ and $50$, contemporaneous consumption falls by 4.67%, 7.89%, and 20.23%, respectively, with effects lasting roughly 2–6 years. This nonlinearity reflects the fact that low-wealth households are constrained in their ability to adjust savings, so consumption responds weakly to preference shocks, whereas wealthy households can reallocate intertemporal resources more flexibly.
@@
The middle row of Figure 3 illustrates heterogeneity with respect to initial risk aversion, holding initial wealth fixed at $w_{t-1}=10$. When initial risk aversion is high ($\gamma_{t-1}=6$), the same shock generates relatively modest responses: consumption falls by 4.84% and the saving rate rises by 5.53%. By contrast, when agents start from a low risk aversion state ($\gamma_{t-1}=1$), consumption drops sharply by 19.26% and saving increases by 32.21%. This asymmetry reflects state-dependent amplification: a positive shock represents a larger proportional change in effective prudence when baseline risk aversion is lower, triggering a stronger reoptimization of intertemporal choices. Consistent with this mechanism, the decline in consumption volatility is also largest at low initial risk aversion.
@@ Appendix A Preliminaries
(19) $(L_{\varphi}^{n}\,h)(z)=\mathbbm{E}_{z}\prod_{t=1}^{n}\varphi_{t}h(Z_{t}),$
where $L_{\varphi}^{n}$ is the $n$-th composition of the operator $L_{\varphi}$ with itself, or equivalently, the $n$-th power of the matrix $L_{\varphi}$. The following result holds. The proof is similar to the proof of Lemma 4.1 of Ma et al. (2020) and thus omitted.